A tree casts a 30 foot shadow. Mi ling, standing next to the tree, casts a 13.5 foot shadow. If mi ling isb5 ft tall, how tall is the tree

Answer by checkley79(3341)

X/30=5/13.5
13.5X=5*30
13.5=150
X=150/13.5
X=11.11 FT. IS THE HEIGHT OF THE TREE,